A proton accelerates from rest in a uniform electric field of640 N/C. At some later time, its speedis 1.30 x 106 m/s.
(a) Find the magnitude of the acceleration ofthe proton.
1m/s2
(b) How long does it take the proton to reach this speed?
2 µs
(c) How far has it moved in that interval?
3 m
(d) What is its kinetic energy at the later time?
4 J
Answer #1
Note: Use the work-kinetic energy theorem (i.e. )
Note: Electric potential is defined as
Note: The initial kinetic energy is zero because the proton wasinitially at REST
Note: The electric force is defined as

Solution (a):








Solution (b):







Solution (c):








Solution (d):
Answer #2
   a.   Acceleration   a   =   F/ m   =   q * E /m   =   1.6 * 10-19 * 640/ 1.67 *10-27
         a   =   6.13* 1010   m/s2
   b.   v   =   u+ a * t
      1.30 *106   =   0   +   6.13* 1010 * t
   time taken to reachspeed   v,         t   =   1.30* 106 / 6.13 *1010   =   2.12 *10-5   s   =   21.2   μs
   c.   distancecovered   d   =   u *t   +   (1/2) * a *t2
         d   =   0* t   +   0.5 * 6.13 *1010 * (2.12 * 10-5)2
         d   =   13.77   m
   d.   k.e.   K   =   (1/2)* m * v2   =   0.5 * 1.67* 10-27 * (1.30 * 106)2
         K   =   1.41* 10-15   J
Answer #3
6.14*10'10M/S
source: homeworklib
answered by: KOACH